OpenStudy (anonymous):

5c^4-3125 is it (c-25)^4 If I am factoring

OpenStudy (anonymous):

No wait it would be 5(c-25)^4

OpenStudy (mathteacher1729):

No, it is neither. \[5c^4-3125 = 5(c^4-625)\] But now we can do something useful with this... Note that 5^4 = 625 so: \[5(c^4-625) = 5(c^4-(5)^4)\] And this is the difference of squares because \[(c^2)^2 = c^4\] So let's use the difference of squares now: \[5(c^4-5^4)= 5(c^2-5^2)(c^2+5^2)\] and again... to give us the final answer! \[5c^4-3125=5(c-5)(c+5)(c^2-5^2)\]

OpenStudy (anonymous):

Thank you I get baffled with this factoring stuff.